A spring is hanging vertically from a fixed point and loaded with a mass of 500 g to its free end, which causes it to stretch by 20.0 cm from its initial length. What is the value of the spring constant?
(-/2.5 Points)
1
Expert's answer
2021-07-12T12:15:34-0400

"\Delta x= 20 cm = 0.2m"

"m = 500g =0.5kg"

"g = 9.8\frac{m}{s^2}"

"\vec F=mg=0.5*9.8=4.9N"

"\vec F=k\Delta x"

"k=\frac{\vec F}{\Delta x}=\frac{4.9}{0.2}=24.5\frac{N}{m}"

"\text{Answer: }k=24.5\frac{N}{m}"